### v1

# HCOA*: Hierarchical Class-ordered A* for Navigation in Semantic Environments

Abstract
This paper addresses the problem of robot navigation in mixed geometric and semantic 3D environments. Given a hierarchical representation of the environment, the objective is to navigate from a start position to a goal while minimizing the computational cost. We introduce Hierarchical Class-ordered A* (HCOA*), an algorithm that leverages the environmental hierarchy for efficient path-planning in semantic graphs, significantly reducing computational effort. We use a total order over the semantic classes and prove theoretical performance guarantees for the algorithm. We propose two approaches for higher-layer node classification based on the node semantics of the lowest layer: a Graph Neural Network-based method and a Majority-Class method. We evaluate our approach through simulations on a 3D Scene Graph (3DSG), comparing it to the state-of-the-art and assessing its performance against our classification approaches. Results show that HCOA* can find the optimal path while reducing the number of expanded nodes by 25% and achieving a 16% reduction in computational time on the uHumans2 3DSG dataset.
I Introduction
As robotic sensing technologies advance, enabling robots to perceive vast and diverse information, two fundamental questions arise: What information from this extensive data stream is most important for a given task, and how can the robot effectively utilize this information for decision-making? Hierarchical semantic environment representations, such as 3D Scene Graphs (3DSG) [1, 2, 3], provide rich and structured abstractions that mirror human way of thinking, facilitating the selection and organization of information.
Previous research in hierarchical path-planning has primarily addressed the first question [4, 5, 6, 7]. In [4] the authors introduce Hierarchical Path-Finding A*, a hierarchical A* variant for grid-based maps. Their approach partitions the map into clusters with designated entrance points, which are used for high-level path-planning. The authors in [5] propose a hierarchical graph search algorithm for graphs with edge weights represented as intervals. Similarly, in [7] the authors propose two hierarchical path-planning methods: path-making and alternating search direction.
Semantic path-planning has focused on the second question by incorporating semantics into the decision-making process. In [8] the authors introduce a weighted function that combines the edge cost and the semantic class to determine the optimal path, while considering the total order of the semantic classes. However, computing this weighted function requires global graph properties, which can be computationally demanding. To address this limitation, [9] and [10] propose Class-ordered A*/LPA*, two extensions of traditional A* [11] that efficiently incorporate semantics, based on a total order over the available classes. Building on the approaches of [5] and [9], in this paper, we develop a hierarchical class-ordered path-planning algorithm, enhancing scalability to larger and more complex environments.
Navigation within hierarchical semantic environments has also been explored in prior work. In [12], the authors introduce D-Lite, a set of two greedy, navigation-oriented compression algorithms for 3DSGs. These algorithms compress the graph based on the shortest path in the lowest layer, and the compressed representation is used for robot navigation. In [13], the authors address Task and Motion Planning in 3DSGs using a three-level hierarchical planner consisting of a discrete task planner, a navigation planner for path-planning, and a low-level planner. Their approach focuses on optimizing task planning to reduce computational cost, while path-planning operates on the unpruned 3DSG.
Learning-based methods have also been explored to facilitate task execution in a 3DSG. In [14], the authors propose HMS, a neural network-based approach for object search in a 3DSG, leveraging the 3DSG’s hierarchy. In [15] and [3], the authors propose the Neural Tree, a Graph Neural Network (GNN) architecture designed for node classification. Although effective in classifying higher-layer nodes in 3DSGs, this approach requires a tree decomposition of the input graph, which increases computational time.
Figure 1: 3D Scene Graph generated from the uHumans2 office scene dataset [16] using Hydra [2]. The graph comprises five layers, as shown in the figure. Room nodes are denoted as R($\cdot$), while building nodes are represented as B($\cdot$).
Main Contributions: Our approach integrates task semantics directly into the path-planning process within a unified algorithm. By leveraging the hierarchical structure of the environment, we significantly reduce computational resource demands while ensuring efficient navigation. Our key contributions are as follows:
  1. 1.
    We introduce Hierarchical Class-ordered A*, a novel hierarchical semantic path-planning algorithm for navigation in hierarchical semantic environments.
  2. 2.
    We propose two methods for node classification on the higher layers of the hierarchy based on underlying semantic classes: a GNN-based method and a majority-class method.
  3. 3.
    We validate our approach on a publicly available 3DSG dataset, demonstrating its effectiveness in real-world scenarios.
II Preliminaries
We model the environment as a hierarchical semantic graph (HSG). Specifically, let $G=(V,E,\mathcal{K})$ be a graph with $n$ layers, where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges, and $\mathcal{K}={1,\ldots K}$ represents a set of semantic classes ordered by decreasing priority. We denote the set of layers as $L$, where $\ell=0$ is the lowest layer and $\ell=n$ is the highest layer (root). Each layer $\ell$ forms a connected weighted subgraph $G^{l}=(V^{\ell},E^{\ell},\mathcal{K})$, with an associated weight function $w:E^{\ell}\to\mathbb{R}^{+}$. We assume that each node in layer $\ell-1$, for $\ell=1,\ldots n$, is connected to exactly one node in layer $\ell$, which we refer to as its parent node. More generally, we define an ancestor as a node’s parent or any higher-layer predecessor in the hierarchy. We define the projection function $p:V\times L\to V$ that maps each node to its corresponding ancestor in layer $\ell$. An example of an HSG is the 3D Scene Graph (3DSG) shown in Figure 1.
Each node in $\ell=0$ is assigned a semantic class based on perception data using the function $\phi_{V}^{0}:V^{0}\to\mathcal{K}$. We extend the labeling function $\phi_{V}^{0}$ to layers $\ell\neq 0$. Details on the computation of $\phi_{V}^{\ell}$ for $\ell\neq 0$ will be given in Section V. Additionally, we introduce the function $\phi_{E}^{\ell}:E^{\ell}\to\mathcal{K}$, which assigns semantic classes to the edges of $G$. Specifically, we define the edge classification function for an edge $e=(v,u)$ as $\phi_{E}^{\ell}(e)=\max(\phi_{V}^{\ell}(v),\phi_{V}^{\ell}(u))$, thereby overestimating the edge class.
Let $\Pi(v_{s}^{\ell},v_{g}^{\ell})$ denote the set of all acyclic paths in $G^{\ell}$ from $v_{s}^{\ell}\in V^{\ell}$ to $v_{g}^{\ell}\in V^{\ell}$, and let $\Pi_{k}$ be the subset of $\Pi$ in which the least favorable (i.e., highest) edge class is exactly $k$. Formally,
$\Pi_{k}=\{\pi:N(\pi,k)>0\text{ and }N(\pi,k^{\prime})=0,\forall k^{\prime}>k\},$ (1)
where $N(\pi,k)=|\{e\in\pi:\phi_{E}^{\ell}(e)=k\}|$ is the number of edges of class $k$ in path $\pi$. We further define $\Pi_{k}^{i}\subseteq\Pi_{k}$ consisting of paths that contain exactly $i$ edges of class $k$, that is $\Pi_{k}^{i}=\{\pi\in\Pi_{k}:N(\pi,k)=i\}$. To enable a consistent comparison of paths across different classes, we impose a total order such that $k<\ell\Rightarrow\Pi_{k}^{i}\prec\Pi_{\ell}^{j}$ for all $\text{for all }i,j$ and $i<j\Rightarrow\Pi_{k}^{i}\prec\Pi_{k}^{j}$. This order ensures that any two paths with the same start and goal nodes can be compared.
Algorithm 1 Hierarchical Class-ordered A* (HCOA*)
Input: $G$, $v_{s}$, $v_{g}$, $h(v)$
      Output: $\pi^{0}$
1:for all $\ell=\ell_{n},\ldots,\ell_{0}$ do
2:     $v_{s}^{\ell}\leftarrow p(v_{s},\ell);\hskip 5.0ptv_{g}^{\ell}\leftarrow p(v_{g},\ell)$
3:     $g(v_{s}^{\ell})\leftarrow 0;\hskip 5.0pt\theta(v_{s}^{\ell})\leftarrow 0\cdot\mathbf{1}_{K};\hskip 5.0ptf(v_{s}^{\ell})\leftarrow h(v_{s}^{\ell})$
4:     $g(v^{\ell})\leftarrow\infty;\hskip 5.0pt\theta(v^{\ell})\leftarrow\infty\cdot\mathbf{1}_{K},\quad\forall v^{\ell}\in G\setminus\{v_{s}^{\ell}\}$
[…]
6:     $Q\leftarrow\{v_{s}^{\ell}\}$triangleright Initialize priority queue
7:     while Q is not empty do
8:         $v\leftarrow$PopNode($Q$);  $Q\leftarrow Q\setminus\{v\}$
9:         if  $v=v_{g}^{\ell}$  then
10:              $\pi^{\ell}\leftarrow$ Path(PredecessorMap, $v_{g}$)
11:              Break
12:         end if
13:         for all $u\in\text{neighbors}(v,G^{\ell^{\prime}})$ do
14:              $\theta(v,u)\leftarrow$ Semantics($\phi_{V}^{\ell}(v),G^{0^{\prime}}(u)$)
15:              if  ($\theta(v)+\theta(v,u)\prec\theta(u)$) or
16:             ($\theta(v)+\theta(v,u)=\theta(u)$ and
17:               $g(v)+w(v,u)<w(u)$) then
18:                  $Q\leftarrow Q\cup\{u\}$
19:                  PredecessorMap[$u$] $\leftarrow v$
20:                  $\theta(u)\leftarrow\theta(v)+\theta(v,u)$
21:                  $g(u)\leftarrow g(v)+w(v,u)$
22:                  $f(u)\leftarrow g(u)+h(u)$
23:              end if
24:         end for
25:     end while
26:     for all $\ell^{\prime}<\ell$ do
27:         $G^{\ell^{\prime}}\leftarrow\{v\in G^{\ell^{\prime}}:p(v,\ell)\in\pi^{\ell}\}$
28:     end for
29:end for
30:return $\pi^{0}$
III Problem Formulation
Consider a mobile robot tasked with navigating a complex 3D environment where certain regions have lower traversal priority. The robot is provided with a HSG of the environment (e.g., 3DSG), as outlined in Section II, where the semantic classes of the nodes in layer $\ell=0$ are assigned based on perception and tailored to the specific task (e.g. avoid certain objects). Let $v_{s}\in V^{0}$ and $v_{g}\in V^{0}$ denote the robot’s starting and goal nodes, respectively. The objective is to determine the shortest path within layer $\ell=0$ while minimizing traversal through the least favorable edges. Formally,
$\pi^{*}(v_{s},v_{g})=\operatorname*{arg\,min}\limits_{\pi\in\Pi^{*}(v_{s},v_{g})}{\sum\limits_{e\in\pi}w(e)},$ (2a)
$\Pi^{*}=\min\limits_{i\in\mathbb{N}}\min\limits_{k\in\mathcal{K}}\Pi_{k}^{i},$ (2b)
where $\Pi^{*}$ is the set of paths obtained by minimizing over all possible classes $k$ and number of least favorable class $i$.
However, due to computational constraints, the robot seeks to avoid a full graph search over the entire $G^{0}$, as its structure is both large and semantically diverse, potentially rendering the search intractable.
III-A Problem Statement
We propose a hierarchical semantic path-planning algorithm for efficient robot navigation in large-scale environments. The algorithm operates top-down across the layers of the HSG, iteratively computing the optimal semantic path at each layer while pruning nodes that are not included in the selected path. Additionally, we introduce two methods for node classification to predict the semantic classes of higher-layer nodes: a Graph Neural Network (GNN)-based approach and a majority-class method.
IV Path-Planning
IV-A Hierarchical Class-ordered A*
We introduce Hierarchical Class-ordered A* (HCOA*), a hierarchical semantic path-planning algorithm for HSGs. The algorithm begins by finding a path in layer $\ell={n-1}$ and then refining it by recursively applying the same process at the lower levels of the hierarchy. By structuring the search hierarchically, the algorithm performs multiple searches on smaller subgraphs rather than executing a computationally expensive search over $G^{0}$. At each layer, the algorithm utilizes Class-ordered A* (COA*) [9], which finds the shortest path while minimizing the number of least favorable edges through lexicographic comparison.
HCOA* is presented in Algorithm 1. We use the function $h:V\to\mathbb{R}$ to denote an admissible heuristic function, similar to standard A*. Lines 2–6 initialize the variables, where the function $p(v,\ell)$ returns the ancestor of node $v$ in layer $\ell$. Lines 7–24 execute COA*, which will be detailed in Section IV-B. Notably, Line 14 computes the semantic class of the edge $(v,u)$ based on the semantic classes of nodes $v$ and $u$. If $\ell=0$, the nodes’ semantics are determined from perception data, whereas for $\ell\neq 0$, they are predicted using the methods described in Section V. Additionally, $G^{0^{\prime}}(u)$ is the induced subgraph of node $u$ in layer $\ell=0$ given by $G^{0^{\prime}}(u)=\{u^{\prime}\in V^{0}:p(u^{\prime},\ell)=u\}$. Finally, Lines 25–27 refine the path by pruning nodes that do not share an ancestor with the paths in the higher layers.
IV-B Class-ordered A*
In this section, we present an overview of the simplified COA* implementation used in HCOA* (Lines 7–24 in Algorithm 1) to help the reader understand the simplifications made on the baseline algorithm. Further details, along with proofs of the COA* optimality and completeness, can be found in [9].
Consider a single-layer, weighted, semantic graph $G^{\ell}$. In this algorithm, each node $v$ is characterized by the triple $(q,g,\theta)$, where $q$ is the predecessor node, $g:V^{\ell}\to\mathbb{R}^{+}$ is the cost-to-come function, and $\theta_{V}:V\to\mathbb{N}^{K}$ specifies the number of edges of each semantic class along the path up to node $v$. We extend $\theta_{V}$ to single edges by defining $\theta_{E}:E^{\ell}\to\mathbb{N}^{K}$ as an one-hot vector with all elements zero except at $\phi_{E}$. The subscript is omitted in Algorithm 1. The function PopNode selects the next node to expand lexicographically, and Path constructs the optimal path by backtracking through parent nodes from the goal. Figure 2 presents the path of COA* in a grid world environment.
Figure 2: (a) Grid environment with three semantic classes: road, grass, river. S denotes the starting cell while G denotes the target; (b) Associated graph of the environment and optimal path produced by Class-ordered A*.
Figure 3: Proposed GNN architecture for node $P\in V^{\ell}$ ($\ell\neq 0$) classification, utilizing the semantic classes (green, blue, red) of nodes in $\ell=0$. The dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ is generated by running COA* on subgraphs of induced nodes. The network consists of two MLPs for pre-processing and post-processing, three GNN layers with skip connections, and an average pooling operator. Training is conducted using cross-entropy loss over the semantic classes.
IV-C Performance Guarantees
The following propositions establish the algorithm’s completeness along with sufficient conditions for its optimality. For both propositions, we assume that the assumptions for completeness and optimality of COA* hold [9]. Proposition 2 relies on two key assumptions. While these assumptions may seem restrictive, they often hold in indoor environments, as discussed in Section VI.
Proposition 1
Let $G$ be an HSG and let nodes $v_{s},v_{g}\in V^{0}$. HCOA* is complete, meaning that it is guaranteed to find a path $\pi^{0}=\pi(v_{s},v_{g})$, whenever one exists.
Proof:
Let $\pi^{0}$ exists. The structure of $G$ ensures that for each layer $\ell\in L$, there exists a corresponding path $\pi^{\ell}=\pi(v_{s}^{\ell},v_{g}^{\ell})$, where $v_{s}^{\ell}=p(v_{s},\ell)$ and $v_{g}^{\ell}=p(v_{g},\ell)$.
At each subgraph $G^{\ell}$, COA* is guaranteed to find the optimal path $\pi^{\ell*}$, if one exists [9]. Suppose, ad absurdum, that the graph pruning performed in Lines 25–27 results in $\Pi(v_{s}^{\ell-1},v_{g}^{\ell-1})=\varnothing$, preventing COA* from finding a solution in layer $\ell-1$. This would imply that $\pi^{\ell*}$ is disconnected, contradicting the completeness of COA*. ∎
Proposition 2
Let $G$ be an HSG with nodes $v_{s},v_{g}\in V^{0}$. Suppose that in layer $\ell=1$, the acyclic path $\pi^{1}(v_{s}^{1},v_{g}^{1})$ is unique, and that for all $u_{s},u_{g}\in V^{0}$ sharing the same parent, $P=p(u_{s},1)=p(u_{g},1)$, every node $u$ along the optimal path $\pi^{*}(u_{s},u_{g})$ satisfies $p(u,1)=P$. Then, HCOA* is guaranteed to find the optimal path $\pi^{*}(v_{s},v_{g})$.
Proof:
Define $G^{0^{\prime}}\subseteq G^{0}$ as the pruned graph after $n$ iterations of HCOA*. Suppose, ad absurdum, that a segment of the optimal path $\pi^{*}(v_{s},v_{g})$ was removed, i.e., there exists a sub-path $\widehat{\pi}\subseteq\pi^{*}(v_{s},v_{g})$ such that $\widehat{\pi}\not\subset G^{0^{\prime}}$. Two cases arise: (i) The parents of $\widehat{\pi}$ in layer $\ell=1$ form a cycle that starts and ends at a parent of a node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$. This contradicts the second assumption, which states that the nodes of the optimal path at $\ell=0$ where the starting and goal node share the same parent, must all have the same parent. (ii) The parents of $\widehat{\pi}$ in layer $\ell=1$ are part of a sub-path that starts at the parent of one node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$ and ends at a different parent of another node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$. This contradicts the first assumption, as it implies the existence of two distinct paths in layer $\ell=1$, violating uniqueness. ∎
V Semantic Class Prediction
Let $\phi_{V}^{\ell}:V^{\ell}\to\mathcal{K}$ be the function that assigns semantic classes to nodes in the layers $\ell\neq 0$. Ideally, given an optimal path $\pi^{*}(v_{s},v_{g})$ computed in $G^{0}$, the semantic class of a node $P\in V^{\ell}$ is determined conservatively by selecting the highest semantic class present in the nodes in $\ell=0$ that both belong to the optimal path and are part of the subgraph $G^{0^{\prime}}(P)=\{u\in V^{0}:p(u,\ell)=P\}$ induced by $P$ in layer $\ell=0$. Formally, this is expressed as:
$\phi_{V}^{\ell}(P)=\max\limits_{u\in\pi^{\prime}(P)}{\phi_{V}^{0}(u)},$ (3a)
$\pi^{\prime}(P)=\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}.$ (3b)
This process follows a bottom-up approach, necessitating the computation of the optimal path in layer $\ell=0$. Given that HCOA* works top-down, we seek an alternative method to assign semantic classes to the higher layers. In the next sections, we present the proposed approaches to predict the semantic classes of higher-layer nodes.
V-A Graph Neural Network Method
We formulate this problem as a supervised node classification task and design a GNN to predict the semantic classes. We construct a dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ by selecting nodes $P\in V^{\ell}$ for $\ell\neq 0$ and extracting their induced subgraphs $G^{0^{\prime}}(P)$. We then assign semantic classes to the nodes in $G^{0^{\prime}}(P)$ based on the given task. For instance, if the task is navigation with object avoidance, we can assign different priorities to regions by placing disks of a specified radius and semantic class around certain object nodes. Next, we execute COA* on this graph and compute $\phi_{V}^{\ell}(P)$, where the starting location is a border node.
A border node $v^{*}\in V^{0}$ of layer $\ell$ is defined as a node that shares an edge with another node from a different ancestor. Formally, there exists $e=(v^{*},u^{*})$ where $v^{*},u^{*}\in V^{0}$ and $p(v^{*},\ell)\neq p(u^{*},\ell)$. Border nodes play a crucial role in the classification of $P\in V^{\ell}$. If the start and goal nodes do not share the same ancestor, the optimal path must traverse a border node in layer $\ell$. Conversely, if both nodes have $P$ as their ancestor, the classification of $P$ becomes irrelevant, as HCOA* will not need to compute a path in layer $\ell$.
The proposed model is illustrated in Figure 3. We use border nodes as input features for the network, which are concatenated with the semantic classes of the nodes in $G^{0^{\prime}}$ represented as one-hot encodings. A 2-layer MLP is used to preprocess the input.
The GNN consists of three layers, each incorporating a Graph Convolutional Network (GCN) operator [17] for message passing, followed by batch normalization and a ReLU activation. To mitigate oversmoothing, we introduce skip connections between the GNN layers. Additionally, we apply average pooling to handle environments of varying sizes. The pooled representation is then processed through another 2-layer MLP, followed by a softmax function. Finally, we use cross-entropy as the loss function for node classification. Formally, given a dataset $\mathcal{D}$ and learnable parameters $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$, the predicted semantic class is obtained by:
$\min\limits_{\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}}{\sum\limits_{(G^{0^{\prime}}(P),\phi_{V}^{\ell}(P))\in\mathcal{D}}{\mathcal{L}_{\text{CE}}\big{(}\widehat{\phi}_{V}^{\ell}(P),\phi_{V}^{\ell}(P)\big{)}}},$ (4)
where $\widehat{\phi}_{V}^{\ell}=f(G^{0^{\prime}}(P);\vartheta_{\text{GNN}},\vartheta_{\text{MLP}})$ represents the model function parametrized by $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$.
V-B Majority-Class Method
The Majority-Class (MC) method is a simple and fast approach for predicting the semantic class of a node $P\in V^{\ell}$ for $\ell\neq 0$ by directly calculating the semantic classes of the nodes in $G^{0^{\prime}}(P)$. Specifically, the predicted class is determined by counting the occurrences of each semantic class among the nodes in the induced subgraph $G^{0^{\prime}}(P)$ and selecting the most frequent one. That is,
$\widehat{\phi}_{V}^{\ell}(P)=\operatorname*{arg\,max}\limits_{k\in\mathcal{K}}{N_{V}(G^{0^{\prime}}(P),k)},$ (5)
where $N_{V}(G,k)=|\{v\in G:\phi_{V}^{0}(v)=k\}|$ is the number of nodes of class $k$ in $G$.
VI Simulations
We perform two sets of simulations on the publicly available uHumans2 office 3DSG, shown in Figure 1 and constructed using Hydra [2]. In the figure, the corresponding layers are also depicted. The places layer $\ell=0$ is a subgraph where each node represents an obstacle-free location, and edges indicate straight-line traversability. The rooms at layer $\ell=1$ consist of nodes representing room centers, with edges connecting neighboring rooms. The number of border nodes (bn) and places in each room are: R(0): 49 bn, 330 places, R(1): 22 bn, 259 places, R(2): 30 bn, 196 places, R(3): 32 bn, 195 places, R(4): 4 bn, 63 places, R(5): 35 bn, 206 places, R(6): 14 bn, 65 places.
We consider a set of three semantic classes $\mathcal{K}$, ordered by decreasing priority: 1 (Green), 2 (Blue), 3 (Red). We utilize the objects layer to identify objects and assign semantic classes to the surrounding place nodes.
In the first section, we present the training results of the GNN and its comparison with the MC approach. The second section demonstrates a path-planning scenario on the entire 3DSG and compares HCOA*, using both the GNN and MC methods, against COA*. HCOA*-GNN utilizes the best GNN model from the previous training phase for room inference. All the simulations were performed using Python 3.8.10 and PyTorch 2.4.1 on a computer with 2.2 GHz, 12-Core, Intel Core i7-8750H CPU, 16GB RAM and an Nvidia RTX 2060 GPU, 6GB VRAM.
VI-A Performance Metrics
To evaluate the performance of room classification, we compute the accuracy, which measures the proportion of correctly classified rooms in each dataset:
$\text{Accuracy}=\frac{\sum\limits_{i=1}^{n}\mathbbm{1}(\widehat{\phi}_{V}^{\ell}(P_{i})=\phi_{V}^{\ell}(P_{i}))}{n},$ (6)
where $n$ is the number of samples, and $\mathbbm{1}(\cdot)$ is the indicator function, which equals 1 if the predicted class matches the ground truth and 0 otherwise.
For the path-planning scenario in the whole 3DSG, we compute the computational time of the algorithms along with the number of expanded nodes (Line 8 in Algorithm 1).
VI-B Room Node Classification
TABLE I: Room Node Classification.
Metrics GNN MC
Training Time (min) 90 -
Validation Accuracy (%) 67.64 30.36
Test Accuracy (Rooms with 1-20 bn) (%) 82.50 54.25
Test Accuracy (Rooms with 21-30 bn) (%) 65.00 19.75
Test Accuracy (Rooms with 31-40 bn) (%) 64.00 19.25
Test Accuracy (Rooms with 41-50 bn) (%) 59.00 24.50
In this section, we present the results from predicting the semantic class of room nodes. To generate the dataset, we constructed 14,000 graphs by extracting the induced subgraphs from all rooms and randomly assigning semantic classes to nodes within a disk of randomly chosen center locations, repeating this process 2,000 times per room. Then, we ran COA* to determine the semantic class of the room, starting from a randomly selected border node. The dataset was split into 80% training, 10% validation, and 10% testing. The model was trained using the Adam optimizer [18] with a learning rate of $10^{-2}$, over 1,600 epochs, a dropout rate of 0.2 and a batch size of 64. The GNN and MLP layers contain 32 neurons.
Table I summarizes the results of room node classification. For the test set, we analyze performance separately for rooms with different numbers of border nodes. The results indicate that the GNN approach outperforms significantly the MC method across both the validation and test sets. Additionally, the training accuracy of the GNN is 67.94%.
VI-C Path-Planning on uHumans2 Office Scene
Figure 4: (a) Places subgraph of the uHumans2 office scene (52m×45m) with the assigned semantic classes (green, blue, red). The starting node is denoted by S, and the goal node by G; (b-d) The path of each algorithm is shown in bold black. Expanded nodes are depicted in regular black, while unexpanded nodes in gray: (b) Path of COA* on the places subgraph; (c) Path of HCOA* on the rooms subgraph; (d) Path of HCOA* on the places subgraph.
[…]
The computed path of COA* in the places layer is shown in Figure 4(4(b)). Figures 4(4(c))-(4(d)) illustrate the results of HCOA* in the rooms and places layers, respectively. The figures also show the expanded nodes for each algorithm. In particular, COA * expanded nodes in two additional rooms compared to HCOA*. The exact number of expanded nodes is provided in Table II, where HCOA* demonstrates a 25% reduction in expanded nodes compared to COA*.
Table II also presents the computational time of the algorithms over 1,000 runs. We observe that HCOA*-MC achieves the best performance in terms of computational efficiency, reducing the execution time by 16% compared to COA*. However, HCOA*-GNN has the highest computational time, as GNN inference can be more time-consuming than graph search in small graphs (the entire places subgraph contains only 1,314 nodes). It should be noted that the first run of HCOA*-GNN takes longer, likely due to compilation and GPU/CPU warm-up overhead. Additionally, both GNN and MC classify R(0) as class 2 and R(2) as class 3. However, while the GNN assigns class 3 to R(3), MC classifies it as class 1. The remaining rooms are not classified, as they are not adjacent to any expanded nodes.
Finally, all three approaches successfully compute the optimal path in the places layer, as shown in Figure 4. It is important to note that while the first assumption of Proposition 2 holds, the second assumption does not. However, this does not affect the results in this case, as the optimal path in the places layer does not require deviations into rooms outside the unique path in the rooms layer to achieve a lower-cost trajectory.
TABLE II: Path-Planning on uHumans2 Office Scene.
Metrics HCOA*-GNN HCOA*-MC COA*
Expanded Nodes 412 412 549
Time ($10^{-3}$s) 38.8 $\pm$ 47.1 4.3 $\boldsymbol{\pm}$ 2.6 5.1 $\pm$ 2.5
Optimal Path
VII Conclusions
In this paper, we addressed the problem of robot navigation in 3D geometric and semantic environments by leveraging the hierarchical structure of the environment for efficient path-planning. We introduced Hierarchical Class-ordered A* (HCOA*), an algorithm that exploits a total order over semantic classes to guide the search process while significantly reducing computational effort. To classify higher-layer nodes, we proposed two methods: a Graph Neural Network-based method and a Majority-Class method. Through simulations on a 3D Scene Graph, we showed that HCOA* effectively reduces computational cost compared to a state-of-the-art method. Specifically, our results show that HCOA* finds the optimal path while reducing the number of expanded nodes by 25% and achieving a 16% reduction in computational time for a typical 3DSG with 1,314 nodes.
Moving forward, future work will focus on applying HCOA* to larger, real-world datasets to demonstrate the full potential of the algorithm. Additionally, we aim to develop an online path-planning framework based on HCOA*, enabling adaptive decision-making in real-time, dynamic scenarios.

### v2

@@
 Abstract
-This paper addresses the problem of robot navigation in mixed geometric and semantic 3D environments.
-Given a hierarchical representation of the environment, the objective is to navigate from a start position to a goal while minimizing the computational cost.
-We introduce Hierarchical Class-ordered A* (HCOA*), an algorithm that leverages the environmental hierarchy for efficient path-planning in semantic graphs, significantly reducing computational effort.
+This paper addresses the problem of robot navigation in mixed geometric/semantic 3D environments.
+Given a hierarchical representation of the environment, the objective is to navigate from a start position to a goal, while satisfying task-specific safety constraints and minimizing computational cost.
+We introduce Hierarchical Class-ordered A* (HCOA*), an algorithm that leverages the environment’s hierarchy for efficient and safe path-planning in mixed geometric/semantic graphs.
 We use a total order over the semantic classes and prove theoretical performance guarantees for the algorithm.
-We propose two approaches for higher-layer node classification based on the node semantics of the lowest layer: a Graph Neural Network-based method and a Majority-Class method.
-We evaluate our approach through simulations on a 3D Scene Graph (3DSG), comparing it to the state-of-the-art and assessing its performance against our classification approaches.
-Results show that HCOA* can find the optimal path while reducing the number of expanded nodes by 25% and achieving a 16% reduction in computational time on the uHumans2 3DSG dataset.
+We propose three approaches for higher-layer node classification based on the semantics of the lowest layer: a Graph Neural Network method, a k-Nearest Neighbors method, and a Majority-Class method.
+We evaluate HCOA* in simulations on two 3D Scene Graphs, comparing it to the state-of-the-art and assessing the performance of each classification approach.
+Results show that HCOA* reduces the computational time of navigation by up to 50%, while maintaining near-optimal performance across a wide range of scenarios.
+Index Terms:
+Autonomous Vehicle Navigation, Motion and Path Planning, AI-Enabled Robotics.
 I Introduction
-As robotic sensing technologies advance, enabling robots to perceive vast and diverse information, two fundamental questions arise: What information from this extensive data stream is most important for a given task, and how can the robot effectively utilize this information for decision-making?
-Hierarchical semantic environment representations, such as 3D Scene Graphs (3DSG) [1, 2, 3], provide rich and structured abstractions that mirror human way of thinking, facilitating the selection and organization of information.
+As robotic sensing technologies advance, enabling robots to perceive vast and diverse information, two fundamental questions arise: What information from this extensive data stream is most important for a given task?
+and, second, how can the robot effectively utilize this information for decision-making?
+Hierarchical semantic environment representations, such as 3D Scene Graphs (3DSGs) [1, 2, 3], provide rich and structured abstractions that mirror human-like reasoning, thus facilitating the selection and organization of information.
 Previous research in hierarchical path-planning has primarily addressed the first question [4, 5, 6, 7].
-In [4] the authors introduce Hierarchical Path-Finding A*, a hierarchical A* variant for grid-based maps.
+In [4] the authors introduce Hierarchical Path-Finding A*, a hierarchical A* [8] variant for grid-based maps.
 Their approach partitions the map into clusters with designated entrance points, which are used for high-level path-planning.
 The authors in [5] propose a hierarchical graph search algorithm for graphs with edge weights represented as intervals.
-Similarly, in [7] the authors propose two hierarchical path-planning methods: path-making and alternating search direction.
-Semantic path-planning has focused on the second question by incorporating semantics into the decision-making process.
-In [8] the authors introduce a weighted function that combines the edge cost and the semantic class to determine the optimal path, while considering the total order of the semantic classes.
-However, computing this weighted function requires global graph properties, which can be computationally demanding.
-To address this limitation, [9] and [10] propose Class-ordered A*/LPA*, two extensions of traditional A* [11] that efficiently incorporate semantics, based on a total order over the available classes.
-Building on the approaches of [5] and [9], in this paper, we develop a hierarchical class-ordered path-planning algorithm, enhancing scalability to larger and more complex environments.
+In [7], the authors introduce S-Nav, a hierarchical path-planning algorithm for navigation in Situational Graphs, a multi-layer graph representation of indoor environments built on top of 3DSGs.
+Semantic path-planning has focused on the second question above by incorporating semantics into the decision-making process.
+In [9], the authors propose SMaNa, a semantic-aware mapping and navigation framework that employs A* with edge weights computed using node distances and semantic labels.
+To avoid reliance on manually crafted, user-defined cost functions, several works adopted the concept of total order over semantic classes.
+In [10], a weighted function combining edge cost and semantic class is introduced to determine optimal paths while respecting a total semantic order.
+However, this approach requires global graph properties, making it computationally expensive.
+To address this limitation, the authors in [11] propose Class-ordered A* (COA*), an extension of traditional A* that incorporates semantic information, based on a total order.
+The algorithm is shown to be both complete and optimal under the defined total order and is validated in various planning scenarios.
+This work was later extended to dynamic environments in [12], where LPA* [13] was adapted for online replanning in dynamic maps with semantic information.
+Building on the ideas introduced in [5] and [11], this paper proposes a hierarchical class-ordered path-planning algorithm that uses a total order on semantic classes.
+The notion of total order is applicable to a wide range of scenarios, including navigation in partially known environments [11] and safety-critical applications.
+In this work, we focus on encoding a set of safety constraints as relationships between semantic categories, which are implicitly enforced during path-planning.
+Figure 1: 3D Scene Graph generated from the uHumans2 office scene dataset [14] using Hydra [2].
+The graph comprises five layers, as shown in the figure.
+Room nodes are denoted as R($\cdot$), while building nodes are represented as B($\cdot$).
 Navigation within hierarchical semantic environments has also been explored in prior work.
-In [12], the authors introduce D-Lite, a set of two greedy, navigation-oriented compression algorithms for 3DSGs.
-These algorithms compress the graph based on the shortest path in the lowest layer, and the compressed representation is used for robot navigation.
-In [13], the authors address Task and Motion Planning in 3DSGs using a three-level hierarchical planner consisting of a discrete task planner, a navigation planner for path-planning, and a low-level planner.
+In [15], the authors present IntelliMove, a hierarchical semantic map framework along with a semantic path-planner.
+The edge weights in the graph encode spatial relationships, such as distances between rooms or objects, and Dijkstra’s algorithm is used to compute optimal paths.
+Furthermore, [16] addresses Task and Motion Planning in 3DSGs using a three-level hierarchical planner consisting of a discrete task planner, a navigation planner for path-planning, and a low-level planner.
 Their approach focuses on optimizing task planning to reduce computational cost, while path-planning operates on the unpruned 3DSG.
@@
 Although effective in classifying higher-layer nodes in 3DSGs, this approach requires a tree decomposition of the input graph, which increases computational time.
-Figure 1: 3D Scene Graph generated from the uHumans2 office scene dataset [16] using Hydra [2].
-The graph comprises five layers, as shown in the figure.
-Room nodes are denoted as R($\cdot$), while building nodes are represented as B($\cdot$).
 Main Contributions: Our approach integrates task semantics directly into the path-planning process within a unified algorithm.
-By leveraging the hierarchical structure of the environment, we significantly reduce computational resource demands while ensuring efficient navigation.
+By leveraging the environment’s hierarchy, we reduce computational resource demands while ensuring efficient and safe navigation.
 Our key contributions are as follows:
@@
   2. 2.
-    We propose two methods for node classification on the higher layers of the hierarchy based on underlying semantic classes: a GNN-based method and a majority-class method.
+    We propose three methods for node classification on the higher layers of the environment’s hierarchy: a Majority-Class (MC) method, a k-Nearest Neighbors (kNN) method, and a GNN method.
   3. 3.
-    We validate our approach on a publicly available 3DSG dataset, demonstrating its effectiveness in real-world scenarios.
+    We validate our approach on two publicly available 3DSG datasets, demonstrating its effectiveness in real-world scenarios. [^1] The code is available at https://github.com/epsomiadis3/HCOA-star.
 II Preliminaries
 We model the environment as a hierarchical semantic graph (HSG).
-Specifically, let $G=(V,E,\mathcal{K})$ be a graph with $n$ layers, where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges, and $\mathcal{K}={1,\ldots K}$ represents a set of semantic classes ordered by decreasing priority.
-We denote the set of layers as $L$, where $\ell=0$ is the lowest layer and $\ell=n$ is the highest layer (root).
-Each layer $\ell$ forms a connected weighted subgraph $G^{l}=(V^{\ell},E^{\ell},\mathcal{K})$, with an associated weight function $w:E^{\ell}\to\mathbb{R}^{+}$.
-We assume that each node in layer $\ell-1$, for $\ell=1,\ldots n$, is connected to exactly one node in layer $\ell$, which we refer to as its parent node.
+Specifically, let $G=(V,E,\mathcal{K})$ be a graph with $n$ layers, where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges, and $\mathcal{K}={1,\ldots K}$ represents a set of semantic classes ordered in decreasing priority.
+This priority is characterized by task-specific requirements (e.g., avoid certain areas or objects, or prefer known over unknown regions).
+We denote the set of layers as $L$, where $\ell=0$ is the lowest abstraction (e.g., sensor) layer and $\ell=n$ is the highest abstraction layer (root).
+Each layer $\ell$ forms a connected weighted subgraph $G^{\ell}=(V^{\ell},E^{\ell},\mathcal{K})\subset G$, with an associated weight function $w^{\ell}:E^{\ell}\to\mathbb{R}^{+}$, which corresponds to Euclidean distance in this work.
+We assume that each node in layer $\ell-1$, for $\ell=1,\ldots n$, is connected to a single node in layer $\ell$, which we refer to as its parent node.
 More generally, we define an ancestor as a node’s parent or any higher-layer predecessor in the hierarchy.
 We define the projection function $p:V\times L\to V$ that maps each node to its corresponding ancestor in layer $\ell$.
-An example of an HSG is the 3D Scene Graph (3DSG) shown in Figure 1.
-Each node in $\ell=0$ is assigned a semantic class based on perception data using the function $\phi_{V}^{0}:V^{0}\to\mathcal{K}$.
+Each node in $\ell=0$ is assigned a semantic class based on the given perception data and the specific task using the function $\phi_{V}^{0}:V^{0}\to\mathcal{K}$.
 We extend the labeling function $\phi_{V}^{0}$ to layers $\ell\neq 0$.
-Details on the computation of $\phi_{V}^{\ell}$ for $\ell\neq 0$ will be given in Section V.
-Additionally, we introduce the function $\phi_{E}^{\ell}:E^{\ell}\to\mathcal{K}$, which assigns semantic classes to the edges of $G$.
-Specifically, we define the edge classification function for an edge $e=(v,u)$ as $\phi_{E}^{\ell}(e)=\max(\phi_{V}^{\ell}(v),\phi_{V}^{\ell}(u))$, thereby overestimating the edge class.
-Let $\Pi(v_{s}^{\ell},v_{g}^{\ell})$ denote the set of all acyclic paths in $G^{\ell}$ from $v_{s}^{\ell}\in V^{\ell}$ to $v_{g}^{\ell}\in V^{\ell}$, and let $\Pi_{k}$ be the subset of $\Pi$ in which the least favorable (i.e., highest) edge class is exactly $k$.
+Details on the computation of $\phi_{V}^{\ell}$ for $\ell\neq 0$ are given in Section V.
+Additionally, we define the edge classification function $\phi_{E}^{\ell}:E^{\ell}\to\mathcal{K}$ by $\phi_{E}^{\ell}(e)=\max(\phi_{V}^{\ell}(v),\phi_{V}^{\ell}(u))$ for $e=(v,u)$, thereby overestimating the edge class.
+The environment model reflects the structure of 3DSGs, as shown in Figure 1.
+For an overview of the advantages of this perceptual model and its construction, we refer the interested reader to [2].
+In our framework, in addition to the spatial layers, we also utilize the object layer of the 3DSG to encode safety constraints and to infer the nodes’ semantic classes, enabling more context-aware and task-relevant navigation.
+Let $\Pi(v_{s}^{\ell},v_{g}^{\ell})$ denote the set of all acyclic paths in $G^{\ell}$ from $v_{s}^{\ell}\in V^{\ell}$ to $v_{g}^{\ell}\in V^{\ell}$, and let $\Pi_{k}$ be the subset of all paths in $\Pi$ for which the least favorable (i.e., highest) edge class is exactly $k$.
 Formally,
-$\Pi_{k}=\{\pi:N(\pi,k)>0\text{ and }N(\pi,k^{\prime})=0,\forall k^{\prime}>k\},$ (1)
+$\Pi_{k}=\{\pi\in\Pi:N(\pi,k)>0;\>N(\pi,k^{\prime})=0,\forall k^{\prime}>k\},$ (1)
 where $N(\pi,k)=|\{e\in\pi:\phi_{E}^{\ell}(e)=k\}|$ is the number of edges of class $k$ in path $\pi$.
-We further define $\Pi_{k}^{i}\subseteq\Pi_{k}$ consisting of paths that contain exactly $i$ edges of class $k$, that is $\Pi_{k}^{i}=\{\pi\in\Pi_{k}:N(\pi,k)=i\}$.
-To enable a consistent comparison of paths across different classes, we impose a total order such that $k<\ell\Rightarrow\Pi_{k}^{i}\prec\Pi_{\ell}^{j}$ for all $\text{for all }i,j$ and $i<j\Rightarrow\Pi_{k}^{i}\prec\Pi_{k}^{j}$.
+We further define $\Pi_{k}^{i}\subseteq\Pi_{k}$ consisting of all paths in $\Pi_{k}$ that contain exactly $i$ edges of class $k$, that is $\Pi_{k}^{i}=\{\pi\in\Pi_{k}:N(\pi,k)=i\}$.
+To enable a consistent comparison of paths across different classes, we impose a total order such that $k<\ell\Rightarrow\Pi_{k}^{i}\prec\Pi_{\ell}^{j}$ $\text{for all }i,j$ and $i<j\Rightarrow\Pi_{k}^{i}\prec\Pi_{k}^{j}$.
 This order ensures that any two paths with the same start and goal nodes can be compared.
 Algorithm 1 Hierarchical Class-ordered A* (HCOA*)
-Input: $G$, $v_{s}$, $v_{g}$, $h(v)$
+Input: $G$, $v_{s}$, $v_{g}$, $h(\cdot)$
    Output: $\pi^{0}$
-1:for all $\ell=\ell_{n},\ldots,\ell_{0}$ do
+1:for all $\ell=\ell_{n-1},\ldots,\ell_{0}$ do
 2:  $v_{s}^{\ell}\leftarrow p(v_{s},\ell);\hskip 5.0ptv_{g}^{\ell}\leftarrow p(v_{g},\ell)$
 3:  $g(v_{s}^{\ell})\leftarrow 0;\hskip 5.0pt\theta(v_{s}^{\ell})\leftarrow 0\cdot\mathbf{1}_{K};\hskip 5.0ptf(v_{s}^{\ell})\leftarrow h(v_{s}^{\ell})$
-4:     $g(v^{\ell})\leftarrow\infty;\hskip 5.0pt\theta(v^{\ell})\leftarrow\infty\cdot\mathbf{1}_{K},\quad\forall v^{\ell}\in G\setminus\{v_{s}^{\ell}\}$
-[…]
-6:     $Q\leftarrow\{v_{s}^{\ell}\}$triangleright Initialize priority queue
+4:  $g(v^{\ell})\leftarrow\infty;\hskip 5.0pt\theta(v^{\ell})\leftarrow\infty\cdot\mathbf{1}_{K},\quad\forall v^{\ell}\in G^{\ell^{\prime}}\setminus\{v_{s}^{\ell}\}$
+5:  PredecessorMap $\leftarrow\varnothing$ $\triangleright$ Track path reconstruction
+6:  $Q\leftarrow\{v_{s}^{\ell}\}$ $\triangleright$ Initialize priority queue
 7:  while Q is not empty do
@@
 13:   for all $u\in\text{neighbors}(v,G^{\ell^{\prime}})$ do
-14:              $\theta(v,u)\leftarrow$ Semantics($\phi_{V}^{\ell}(v),G^{0^{\prime}}(u)$)
+[…]
 15:     if  ($\theta(v)+\theta(v,u)\prec\theta(u)$) or
 16:     ($\theta(v)+\theta(v,u)=\theta(u)$ and
-17:               $g(v)+w(v,u)<w(u)$) then
+17:      $g(v)+w^{\ell}(v,u)<g(u)$) then
 18:      $Q\leftarrow Q\cup\{u\}$
@@
 20:      $\theta(u)\leftarrow\theta(v)+\theta(v,u)$
-21:                  $g(u)\leftarrow g(v)+w(v,u)$
+21:      $g(u)\leftarrow g(v)+w^{\ell}(v,u)$
 22:      $f(u)\leftarrow g(u)+h(u)$
@@
 25:  end while
-26:     for all $\ell^{\prime}<\ell$ do
-27:         $G^{\ell^{\prime}}\leftarrow\{v\in G^{\ell^{\prime}}:p(v,\ell)\in\pi^{\ell}\}$
+26:  for all $\hat{\ell}$$\ell$ do
+27:   $G^{\hat{\ell}^{\prime}}\leftarrow\{v\in G^{\hat{\ell}^{\prime}}:p(v,\ell)\in\pi^{\ell}\}$
 28:  end for
@@
 Consider a mobile robot tasked with navigating a complex 3D environment where certain regions have lower traversal priority.
-The robot is provided with a HSG of the environment (e.g., 3DSG), as outlined in Section II, where the semantic classes of the nodes in layer $\ell=0$ are assigned based on perception and tailored to the specific task (e.g.
-avoid certain objects).
+The robot is provided with a HSG of the environment (e.g., 3DSG), as outlined in Section II, where the semantic classes of the nodes in layer $\ell=0$ are assigned based on perception, and tailored to the specific task (e.g., safety constraints).
 Let $v_{s}\in V^{0}$ and $v_{g}\in V^{0}$ denote the robot’s starting and goal nodes, respectively.
-The objective is to determine the shortest path within layer $\ell=0$ while minimizing traversal through the least favorable edges.
+The objective is to determine the shortest path in $\ell=0$ while minimizing traversal through the least favorable edges.
 Formally,
-$\pi^{*}(v_{s},v_{g})=\operatorname*{arg\,min}\limits_{\pi\in\Pi^{*}(v_{s},v_{g})}{\sum\limits_{e\in\pi}w(e)},$ (2a)
+$\pi^{*}(v_{s},v_{g})=\operatorname*{arg\,min}\limits_{\pi\in\Pi^{*}(v_{s},v_{g})}{\sum\limits_{e\in\pi}w^{0}(e)},$ (2a)
 $\Pi^{*}=\min\limits_{i\in\mathbb{N}}\min\limits_{k\in\mathcal{K}}\Pi_{k}^{i},$ (2b)
@@
 We propose a hierarchical semantic path-planning algorithm for efficient robot navigation in large-scale environments.
-The algorithm operates top-down across the layers of the HSG, iteratively computing the optimal semantic path at each layer while pruning nodes that are not included in the selected path.
-Additionally, we introduce two methods for node classification to predict the semantic classes of higher-layer nodes: a Graph Neural Network (GNN)-based approach and a majority-class method.
+The algorithm operates top-down across the layers of the HSG, iteratively computing the optimal semantic path at each layer while pruning nodes that are not included in the path.
+Additionally, we introduce three methods for node classification to predict the semantic classes of higher-layer nodes: a Majority-Class, a kNN, and a GNN method.
 IV Path-Planning
@@
 We introduce Hierarchical Class-ordered A* (HCOA*), a hierarchical semantic path-planning algorithm for HSGs.
-The algorithm begins by finding a path in layer $\ell={n-1}$ and then refining it by recursively applying the same process at the lower levels of the hierarchy.
-By structuring the search hierarchically, the algorithm performs multiple searches on smaller subgraphs rather than executing a computationally expensive search over $G^{0}$.
-At each layer, the algorithm utilizes Class-ordered A* (COA*) [9], which finds the shortest path while minimizing the number of least favorable edges through lexicographic comparison.
+The algorithm initiates planning at layer $\ell=n-1$, and recursively proceeds to lower layers.
+At each layer, it prunes the HSG based on the computed path from the previous layer and applies the same planning algorithm to the pruned HSG.
+Within each layer, the algorithm utilizes Class-ordered A* (COA*) [11], which finds the shortest path while minimizing the number of least favorable edges through lexicographic comparison.
+By leveraging the hierarchy, the algorithm performs multiple searches on smaller subgraphs rather than a potentially computationally expensive search over $G^{0}$.
 HCOA* is presented in Algorithm 1.
@@
 Notably, Line 14 computes the semantic class of the edge $(v,u)$ based on the semantic classes of nodes $v$ and $u$.
-If $\ell=0$, the nodes’ semantics are determined from perception data, whereas for $\ell\neq 0$, they are predicted using the methods described in Section V.
-Additionally, $G^{0^{\prime}}(u)$ is the induced subgraph of node $u$ in layer $\ell=0$ given by $G^{0^{\prime}}(u)=\{u^{\prime}\in V^{0}:p(u^{\prime},\ell)=u\}$.
-Finally, Lines 25–27 refine the path by pruning nodes that do not share an ancestor with the paths in the higher layers.
+If $\ell=0$, the nodes’ semantics are determined from perception data and the task, whereas for $\ell\neq 0$, they are predicted using the methods described in Section V.
+Additionally, $G^{0^{\prime}}(u)$ is the induced subgraph of node $u\in V^{\ell}$ on layer 0 given by $G^{0^{\prime}}(u)=\{u^{\prime}\in V^{0}:p(u^{\prime},\ell)=u,\>\text{where}\>u\in V^{\ell}\}$, while $G^{\ell^{\prime}}$ is the pruned graph at layer $\ell$ (Line 27).
+Lines 25–27 refine the HSG by pruning nodes that do not share an ancestor with the paths in the higher layers.
 IV-B Class-ordered A*
-In this section, we present an overview of the simplified COA* implementation used in HCOA* (Lines 7–24 in Algorithm 1) to help the reader understand the simplifications made on the baseline algorithm.
+In this section, we present an overview of the simplified COA* implementation used in HCOA* to help the reader understand the simplifications made to the baseline algorithm.
 Further details, along with proofs of the COA* optimality and completeness, can be found in [11].
 Consider a single-layer, weighted, semantic graph $G^{\ell}$.
-In this algorithm, each node $v$ is characterized by the triple $(q,g,\theta)$, where $q$ is the predecessor node, $g:V^{\ell}\to\mathbb{R}^{+}$ is the cost-to-come function, and $\theta_{V}:V\to\mathbb{N}^{K}$ specifies the number of edges of each semantic class along the path up to node $v$.
+We characterize each node $v$ by the triple $(q,g,\theta_{V})$, where $q$ is the predecessor node (saved in the PredecessorMap in Algorithm 1), $g:V^{\ell}\to\mathbb{R}^{+}$ is the cost-to-come function, and $\theta_{V}:V^{\ell}\to\mathbb{N}^{K}$ specifies the number of edges of each semantic class along the path up to node $v$.
 We extend $\theta_{V}$ to single edges by defining $\theta_{E}:E^{\ell}\to\mathbb{N}^{K}$ as an one-hot vector with all elements zero except at $\phi_{E}$.
 The subscript is omitted in Algorithm 1.
-The function PopNode selects the next node to expand lexicographically, and Path constructs the optimal path by backtracking through parent nodes from the goal.
-Figure 2 presents the path of COA* in a grid world environment.
-Figure 2: (a) Grid environment with three semantic classes: road, grass, river.
-S denotes the starting cell while G denotes the target; (b) Associated graph of the environment and optimal path produced by Class-ordered A*.
-Figure 3: Proposed GNN architecture for node $P\in V^{\ell}$ ($\ell\neq 0$) classification, utilizing the semantic classes (green, blue, red) of nodes in $\ell=0$.
-The dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ is generated by running COA* on subgraphs of induced nodes.
-The network consists of two MLPs for pre-processing and post-processing, three GNN layers with skip connections, and an average pooling operator.
-Training is conducted using cross-entropy loss over the semantic classes.
+The function PopNode selects the next node to expand, and Path constructs the optimal path by backtracking through the predecessor nodes from the goal.
+Figure 2: (a) Grid world with three semantic classes: road, grass, river. S is the starting cell while G is the target; (b) Graph of the environment and path produced by COA*.
+Figure 2 presents an example of COA* in a grid world, where the robot must navigate from $\mathbf{S}$ to $\mathbf{G}$ while satisfying safety constraints prioritized in descending order: (i) avoid water, and (ii) avoid grass.
+To enforce these constraints, we assign semantic labels to nodes via $\phi_{V}^{\ell}$.
 IV-C Performance Guarantees
-The following propositions establish the algorithm’s completeness along with sufficient conditions for its optimality.
-For both propositions, we assume that the assumptions for completeness and optimality of COA* hold [9].
-Proposition 2 relies on two key assumptions.
-While these assumptions may seem restrictive, they often hold in indoor environments, as discussed in Section VI.
-Proposition 1
+The following propositions establish the algorithm’s completeness along with sufficient conditions for optimality.
+For both propositions, the assumptions for completeness and optimality of COA* hold [11].
+We conclude the section with a discussion on the algorithm’s computational complexity.
+Proposition 1.
 Let $G$ be an HSG and let nodes $v_{s},v_{g}\in V^{0}$. HCOA* is complete, meaning that it is guaranteed to find a path $\pi^{0}=\pi(v_{s},v_{g})$, whenever one exists.
-Proof:
-Let $\pi^{0}$ exists.
+Proof.
+Let $\pi^{0}$ exist.
 The structure of $G$ ensures that for each layer $\ell\in L$, there exists a corresponding path $\pi^{\ell}=\pi(v_{s}^{\ell},v_{g}^{\ell})$, where $v_{s}^{\ell}=p(v_{s},\ell)$ and $v_{g}^{\ell}=p(v_{g},\ell)$.
@@
 Suppose, ad absurdum, that the graph pruning performed in Lines 25–27 results in $\Pi(v_{s}^{\ell-1},v_{g}^{\ell-1})=\varnothing$, preventing COA* from finding a solution in layer $\ell-1$.
-This would imply that $\pi^{\ell*}$ is disconnected, contradicting the completeness of COA*.
+This implies that $\pi^{\ell*}$ is disconnected, contradicting COA*’s completeness.
 ∎
-Proposition 2
+Proposition 2.
 Let $G$ be an HSG with nodes $v_{s},v_{g}\in V^{0}$.
@@
 Then, HCOA* is guaranteed to find the optimal path $\pi^{*}(v_{s},v_{g})$.
-Proof:
+Proof.
 Define $G^{0^{\prime}}\subseteq G^{0}$ as the pruned graph after $n$ iterations of HCOA*.
-Suppose, ad absurdum, that a segment of the optimal path $\pi^{*}(v_{s},v_{g})$ was removed, i.e., there exists a sub-path $\widehat{\pi}\subseteq\pi^{*}(v_{s},v_{g})$ such that $\widehat{\pi}\not\subset G^{0^{\prime}}$.
+Suppose, ad absurdum, that a segment of the optimal path $\pi^{*}(v_{s},v_{g})$ has been removed, i.e., there exists a sub-path $\widehat{\pi}\subseteq\pi^{*}(v_{s},v_{g})$ such that $\widehat{\pi}\not\subset G^{0^{\prime}}$.
 Two cases arise: (i) The parents of $\widehat{\pi}$ in layer $\ell=1$ form a cycle that starts and ends at a parent of a node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$.
-This contradicts the second assumption, which states that the nodes of the optimal path at $\ell=0$ where the starting and goal node share the same parent, must all have the same parent.
+This contradicts the second assumption, which states that the nodes of the optimal path at $\ell=0$, where the starting and goal node share the same parent, must all have the same parent.
 (ii) The parents of $\widehat{\pi}$ in layer $\ell=1$ are part of a sub-path that starts at the parent of one node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$ and ends at a different parent of another node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$.
@@
 ∎
+The two assumptions of Proposition 2 may seem restrictive at first, but they often hold in indoor environments.
+Figure 3 shows an example, where $\mathbf{G}$ denotes the goal, $R(\cdot)$ denotes a room, and $C_{1}$, $C_{2}$ represent specific conditions, namely $C_{1}$: open door and $C_{2}$: dangerous area.
+When both $C_{1}$ and $C_{2}$ are false, Proposition 2 holds, and HCOA* successfully finds the optimal path (shown in green).
+However, when $C_{1}$ is true, the path produced by HCOA* (still shown in green) becomes suboptimal.
+This occurs because the room is represented, in this example, by its center node.
+For large rooms such as corridors (e.g., $R(1)$), this representation becomes misleading.
+Additionally, if $C_{2}$ is true, the optimal path requires moving from $R(2)$ to $R(1)$ and then back to $R(2)$.
+This cycle cannot be captured by COA* operating at the room layer, as it computes acyclic paths.
+Hence, HCOA* produces a suboptimal path in the places layer in this case (shown in blue).
+In general, the spatial representation of nodes in higher layers, as well as their computed semantic classes, can significantly affect HCOA*’s performance, potentially leading to suboptimal solutions.
+Nevertheless, as demonstrated in Section VI, HCOA* manages to find the optimal path in most tested scenarios.
+One potential improvement is to refine the representation of rooms by associating multiple nodes with each room, for example, using nodes near doors rather than using a single node at the geometric center of the room.
+Figure 3: Example demonstrating Proposition 2.
+$\mathbf{G}$ denotes the goal location, $R(\cdot)$ indicates rooms, and $C_{1}$, $C_{2}$ represent specific conditions: $C_{1}$: open door, and $C_{2}$: dangerous area.
+Figure 4: Proposed GNN architecture for node $P\in V^{\ell}$ ($\ell\neq 0$) classification, utilizing the semantic classes (green, blue, red) of nodes in $\ell=0$.
+The dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ is generated by running COA* on induced subgraphs of nodes.
+The network consists of two MLPs for pre-processing and post-processing, three GNN layers with skip connections, and an average pooling operator.
+Training is conducted using cross-entropy loss over the semantic classes.
+Regarding the computational complexity of the algorithm, note that, in the worst case, COA* with an admissible heuristic has the same complexity as Dijkstra’s algorithm, which is $\mathcal{O}(|E^{0}|+|V^{0}|\log|V^{0}|)$ when implemented with a Fibonacci heap.
+Hence, the worst-case complexity of HCOA* is given by $\mathcal{O}\left(\sum_{\ell=0}^{L}\left(|E^{\ell^{\prime}}|+|V^{\ell^{\prime}}|\log|V^{\ell^{\prime}}|\right)\right)$, where $|V^{\ell}{}^{\prime}|$ and $|E^{\ell}{}^{\prime}|$ denote the number of vertices and edges, respectively, of the pruned graph at each layer $\ell$.
 V Semantic Class Prediction
-Let $\phi_{V}^{\ell}:V^{\ell}\to\mathcal{K}$ be the function that assigns semantic classes to nodes in the layers $\ell\neq 0$.
-Ideally, given an optimal path $\pi^{*}(v_{s},v_{g})$ computed in $G^{0}$, the semantic class of a node $P\in V^{\ell}$ is determined conservatively by selecting the highest semantic class present in the nodes in $\ell=0$ that both belong to the optimal path and are part of the subgraph $G^{0^{\prime}}(P)=\{u\in V^{0}:p(u,\ell)=P\}$ induced by $P$ in layer $\ell=0$.
-Formally, this is expressed as:
+Let $\phi_{V}^{\ell}:V^{\ell}\to\mathcal{K}$ be the function that assigns semantic classes to nodes in layers $\ell=1,\ldots,n$.
+Ideally, given an optimal path $\pi^{*}(v_{s},v_{g})$ computed in $G^{0}$, the semantic class of a node $P\in V^{\ell}$ ($\ell=1,\ldots,n$) is determined conservatively by selecting the highest semantic class present in the nodes in $\ell=0$ that both belong to the optimal path and are part of the subgraph $G^{0^{\prime}}(P)=\{u\in V^{0}:p(u,\ell)=P,\>\text{where}\>P\in V^{\ell}\}$ induced by $P$ on layer 0.
+Formally,
 $\phi_{V}^{\ell}(P)=\max\limits_{u\in\pi^{\prime}(P)}{\phi_{V}^{0}(u)},$ (3a)
-$\pi^{\prime}(P)=\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}.$ (3b)
+$\pi^{\prime}(P)=\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}(P).$ (3b)
 This process follows a bottom-up approach, necessitating the computation of the optimal path in layer $\ell=0$.
-Given that HCOA* works top-down, we seek an alternative method to assign semantic classes to the higher layers.
-In the next sections, we present the proposed approaches to predict the semantic classes of higher-layer nodes.
-V-A Graph Neural Network Method
-We formulate this problem as a supervised node classification task and design a GNN to predict the semantic classes.
-We construct a dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ by selecting nodes $P\in V^{\ell}$ for $\ell\neq 0$ and extracting their induced subgraphs $G^{0^{\prime}}(P)$.
+Given that HCOA* works top-down, we seek alternative methods to find the semantic classes of the higher layers.
+We introduce three methods for predicting the semantic class of higher-layer nodes: a Majority-Class (MC), a k-Nearest Neighbors (kNN), and a Graph Neural Network (GNN) approach.
+The latter two methods require a supervised training phase, which necessitates the construction of a dataset.
+To this end, we construct a dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ by selecting nodes $P\in V^{\ell}$ for $\ell\neq 0$ and extracting their induced subgraphs $G^{0^{\prime}}(P)$.
 We then assign semantic classes to the nodes in $G^{0^{\prime}}(P)$ based on the given task.
-For instance, if the task is navigation with object avoidance, we can assign different priorities to regions by placing disks of a specified radius and semantic class around certain object nodes.
-Next, we execute COA* on this graph and compute $\phi_{V}^{\ell}(P)$, where the starting location is a border node.
-A border node $v^{*}\in V^{0}$ of layer $\ell$ is defined as a node that shares an edge with another node from a different ancestor.
+For navigation tasks involving object avoidance, we determine region priorities by placing disks of a specified radius and semantic class around certain objects.
+Next, we run COA* on $G^{0^{\prime}}(P)$ and compute $\phi_{V}^{\ell}(P)$, where the start is a border node.
+A border node $v^{*}\in V^{0}$ is a node that is connected by an edge to another (border) node $u^{*}\in V^{0}$ whose ancestor in layer $\ell$ is different from that of $v^{*}$.
 Formally, there exists $e=(v^{*},u^{*})$ where $v^{*},u^{*}\in V^{0}$ and $p(v^{*},\ell)\neq p(u^{*},\ell)$.
@@
 Conversely, if both nodes have $P$ as their ancestor, the classification of $P$ becomes irrelevant, as HCOA* will not need to compute a path in layer $\ell$.
-The proposed model is illustrated in Figure 3.
-We use border nodes as input features for the network, which are concatenated with the semantic classes of the nodes in $G^{0^{\prime}}$ represented as one-hot encodings.
+V-A Majority-Class Method
+The Majority-Class (MC) method is a simple and fast approach for predicting the semantic class of a node $P\in V^{\ell}$ for $\ell\neq 0$.
+The predicted class is determined by counting the occurrences of each semantic class in the induced subgraph $G^{0^{\prime}}(P)$ and selecting the most frequent one.
+That is,
+$\widehat{\phi}_{V}^{\ell}(P)=\operatorname*{arg\,max}\limits_{k\in\mathcal{K}}{N_{V}(G^{0^{\prime}}(P),k)},$ (4)
+where $N_{V}(G,k)=|\{v\in G:\phi_{V}^{0}(v)=k\}|$ is the number of nodes of class $k$ in $G$.
+V-B k-Nearest Neighbors Method
+The kNN algorithm [19] is a non-parametric, instance-based learning method that classifies data points based on the majority vote of their $\mathsf{k}$ nearest neighbors in the feature space.
+Given the induced subgraph $G^{0^{\prime}}(P)$ with node set $V^{0^{\prime}}$, we extract a graph-level feature vector $\mathsf{f}$ consisting of: (i) the proportion of each semantic class in the graph
+$\frac{N_{V}(G^{0^{\prime}}(P),k)}{|V^{0^{\prime}}|},\quad k\in\mathcal{K},$ (5)
+and (ii) the majority class among the border nodes $\operatorname*{arg\,max}_{k\in\mathcal{K}}{N_{V^{*}}(G^{0^{\prime}}(P),k)}$, where $V^{*}$ denotes the set of border nodes and $k$ the index of a semantic class (not to be confused with $\mathsf{k}$ used in kNN).
+We intentionally limit the number of features, as the inference time of kNN increases with the dimensionality of the feature space.
+To ensure uniform scaling, we standardize all feature vectors $\mathsf{f}$.
+Graph similarity is measured via Euclidean distance, and classification is performed by assigning the majority class among the $\mathsf{k}$ nearest neighbors from the training set:
+$\widehat{\phi}_{V}^{\ell}(P)=\operatorname*{arg\,max}\limits_{k\in\mathcal{K}}\sum\limits_{i=1}^{\mathsf{k}}\mathbbm{1}[\phi_{V}^{\ell}(P_{i})=k],$ (6)
+where $P_{1},P_{2},\ldots,P_{\mathsf{k}}$ are the $\mathsf{k}$ nearest neighbors of $P$ in the training set, and $\mathbbm{1}[\cdot]$ is the indicator function that equals 1 if the condition is true and 0 otherwise.
+V-C Graph Neural Network Method
+The GNN method is a supervised learning approach that leverages the graph structure and features through message passing.
+The proposed model is shown in Figure 4.
+We use border nodes as input features, which are concatenated with the semantic classes of the nodes in $G^{0^{\prime}}$ represented as one-hot encodings.
 A 2-layer MLP is used to preprocess the input.
 The GNN consists of three layers, each incorporating a Graph Convolutional Network (GCN) operator [20] for message passing, followed by batch normalization and a ReLU activation.
+We adopt a 3-layer architecture, as deeper GNNs tend to suffer from oversmoothing and slower inference, both of which are critical limitations in path-planning applications.
 To mitigate oversmoothing, we introduce skip connections between the GNN layers.
@@
 Finally, we use cross-entropy as the loss function for node classification.
-Formally, given a dataset $\mathcal{D}$ and learnable parameters $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$, the predicted semantic class is obtained by:
+Formally, given a dataset $\mathcal{D}$ and learnable parameters $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$, the predicted semantic class is obtained by
 $\min\limits_{\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}}{\sum\limits_{(G^{0^{\prime}}(P),\phi_{V}^{\ell}(P))\in\mathcal{D}}{\mathcal{L}_{\text{CE}}\big{(}\widehat{\phi}_{V}^{\ell}(P),\phi_{V}^{\ell}(P)\big{)}}},$ (7)
-where $\widehat{\phi}_{V}^{\ell}=f(G^{0^{\prime}}(P);\vartheta_{\text{GNN}},\vartheta_{\text{MLP}})$ represents the model function parametrized by $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$.
-V-B Majority-Class Method
-The Majority-Class (MC) method is a simple and fast approach for predicting the semantic class of a node $P\in V^{\ell}$ for $\ell\neq 0$ by directly calculating the semantic classes of the nodes in $G^{0^{\prime}}(P)$.
-Specifically, the predicted class is determined by counting the occurrences of each semantic class among the nodes in the induced subgraph $G^{0^{\prime}}(P)$ and selecting the most frequent one.
-That is,
-$\widehat{\phi}_{V}^{\ell}(P)=\operatorname*{arg\,max}\limits_{k\in\mathcal{K}}{N_{V}(G^{0^{\prime}}(P),k)},$ (5)
-where $N_{V}(G,k)=|\{v\in G:\phi_{V}^{0}(v)=k\}|$ is the number of nodes of class $k$ in $G$.
+where $\widehat{\phi}_{V}^{\ell}(P)=f(G^{0^{\prime}}(P);\vartheta_{\text{GNN}},\vartheta_{\text{MLP}})$ represents the model function parametrized by $\vartheta_{\text{GNN}},\vartheta_{\text{MLP}}$.
 VI Simulations
-We perform two sets of simulations on the publicly available uHumans2 office 3DSG, shown in Figure 1 and constructed using Hydra [2].
-In the figure, the corresponding layers are also depicted.
-The places layer $\ell=0$ is a subgraph where each node represents an obstacle-free location, and edges indicate straight-line traversability.
-The rooms at layer $\ell=1$ consist of nodes representing room centers, with edges connecting neighboring rooms.
-The number of border nodes (bn) and places in each room are: R(0): 49 bn, 330 places, R(1): 22 bn, 259 places, R(2): 30 bn, 196 places, R(3): 32 bn, 195 places, R(4): 4 bn, 63 places, R(5): 35 bn, 206 places, R(6): 14 bn, 65 places.
-We consider a set of three semantic classes $\mathcal{K}$, ordered by decreasing priority: 1 (Green), 2 (Blue), 3 (Red).
-We utilize the objects layer to identify objects and assign semantic classes to the surrounding place nodes.
-In the first section, we present the training results of the GNN and its comparison with the MC approach.
-The second section demonstrates a path-planning scenario on the entire 3DSG and compares HCOA*, using both the GNN and MC methods, against COA*.
-HCOA*-GNN utilizes the best GNN model from the previous training phase for room inference.
+We perform simulations on two publicly available datasets generated with Hydra [2]: the uHumans2 office 3DSG (Figure 1) and the subway 3DSG (Figure 6(6(a))).
+Figure 1 also shows the layer names.
+The place layer ($\ell=0$) is a graph with nodes as obstacle-free locations and edges indicating traversability.
+The room layer ($\ell=1$) consists of nodes representing room centers and edges connecting neighboring rooms.
+Our task involves navigation with object or room avoidance.
+The pipeline proceeds as follows: we first translate safety-related commands (e.g., “avoid computers”) into semantic classes at the place layer, and impose a total order based on their priority.
+Next, depending on the room classification method employed, we train the corresponding models to predict semantic room classes.
+Finally, we select the best-performing model and execute HCOA* to compute the path.
+In all simulations, we impose two safety constraints resulting in three semantic classes $\mathcal{K}$, ordered by decreasing priority.
+We leverage the room and object layers to identify relevant rooms and objects, and assign semantic labels to nearby place nodes accordingly.
+The resulting classes are visualized (in decreasing priority) as: 1 (Green), 2 (Blue), and 3 (Red).
+The first section presents the comparison of the three methods for semantic class prediction.
+The second section showcases path-planning scenarios in the office 3DSG, comparing HCOA* to COA*.
+We use the same weight function across all graph layers and denote it by $w$ for simplicity, corresponding to the Euclidean distance between nodes.
+We also include results from a modified A* algorithm, denoted MA* (similar to [9]), to demonstrate the importance of imposing a total order over semantic classes.
+In this variant, the edge cost in A* is modified to $w^{\prime}=w+\alpha^{k}$, where $k\in\mathcal{K}$ and $\alpha$ is a scaling factor.
+The third section presents analogous experiments in the subway 3DSG, also comparing HCOA* to COA* and MA*.
+In these experiments, HCOA*-GNN utilizes the best GNN model from the previous section for room inference.
 All the simulations were performed using Python 3.8.10 and PyTorch 2.4.1 on a computer with 2.2 GHz, 12-Core, Intel Core i7-8750H CPU, 16GB RAM and an Nvidia RTX 2060 GPU, 6GB VRAM.
 VI-A Performance Metrics
-To evaluate the performance of room classification, we compute the accuracy, which measures the proportion of correctly classified rooms in each dataset:
-$\text{Accuracy}=\frac{\sum\limits_{i=1}^{n}\mathbbm{1}(\widehat{\phi}_{V}^{\ell}(P_{i})=\phi_{V}^{\ell}(P_{i}))}{n},$ (6)
-where $n$ is the number of samples, and $\mathbbm{1}(\cdot)$ is the indicator function, which equals 1 if the predicted class matches the ground truth and 0 otherwise.
-For the path-planning scenario in the whole 3DSG, we compute the computational time of the algorithms along with the number of expanded nodes (Line 8 in Algorithm 1).
-VI-B Room Node Classification
-TABLE I: Room Node Classification.
-Metrics GNN MC
-Training Time (min) 90 -
-Validation Accuracy (%) 67.64 30.36
-Test Accuracy (Rooms with 1-20 bn) (%) 82.50 54.25
-Test Accuracy (Rooms with 21-30 bn) (%) 65.00 19.75
-Test Accuracy (Rooms with 31-40 bn) (%) 64.00 19.25
-Test Accuracy (Rooms with 41-50 bn) (%) 59.00 24.50
+To evaluate the performance of the prediction of a room’s semantic class, we compute the accuracy, which measures the proportion of correctly classified rooms in each dataset:
+$\text{Accuracy}=\frac{\sum\limits_{i=1}^{n}\mathbbm{1}[\widehat{\phi}_{V}^{\ell}(P_{i})=\phi_{V}^{\ell}(P_{i})]}{n},$ (8)
+where $n$ is the number of samples. For the planning scenarios, we compute the algorithms’ computational time along with the number of expanded nodes (Line 8 in Algorithm 1).
+VI-B Semantic Class Prediction
+TABLE I: Semantic Class Prediction.
+Metrics MC kNN GNN
+Training Time - 10 s 97 min
+Validation Acc. (%) 42.79 52.43 63.43
+Test Acc. (1-20 bn Rooms) (%) 57.00 70.50 74.00
+Test Acc. (21-30 bn Rooms) (%) 33.75 48.50 58.25
+Test Acc. (31-40 bn Rooms) (%) 37.00 44.25 54.75
+Test Acc. (41-50 bn Rooms) (%) 41.50 43.00 47.50
 In this section, we present the results from predicting the semantic class of room nodes.
-To generate the dataset, we constructed 14,000 graphs by extracting the induced subgraphs from all rooms and randomly assigning semantic classes to nodes within a disk of randomly chosen center locations, repeating this process 2,000 times per room.
+To generate the dataset, we constructed 14,000 graphs by extracting the induced subgraphs from all rooms in the Office 3DSG and randomly assigning semantic classes to nodes within a disk of randomly chosen center locations, repeating this process 2,000 times per room.
 Then, we ran COA* to determine the semantic class of the room, starting from a randomly selected border node.
 The dataset was split into 80% training, 10% validation, and 10% testing.
-The model was trained using the Adam optimizer [18] with a learning rate of $10^{-2}$, over 1,600 epochs, a dropout rate of 0.2 and a batch size of 64.
+The GNN model was trained using the Adam optimizer [21] with a learning rate of $10^{-2}$, over 1,600 epochs, a dropout rate of 0.2 and a batch size of 64.
 The GNN and MLP layers contain 32 neurons.
+For training the kNN model, we set $\mathsf{k}=5$, and used half of the training dataset.
 Table I summarizes the results of room node classification.
-For the test set, we analyze performance separately for rooms with different numbers of border nodes.
-The results indicate that the GNN approach outperforms significantly the MC method across both the validation and test sets.
-Additionally, the training accuracy of the GNN is 67.94%.
+For the test set, we analyzed the performance separately for rooms with different numbers of border nodes (bn).
+The results indicate that the GNN approach outperforms the other two methods.
+This result is expected, as the GNN learns complex spatial, structural, and semantic features, unlike the MC baseline and the kNN method, which relies on hand-crafted features.
+Furthermore, we observe that all methods achieve higher accuracy for rooms with fewer bns.
+These rooms are typically smaller and have only one entrance, resulting in spatially clustered border nodes.
+In such cases, features considering the border nodes’ semantics lead to improved classification performance.
+The training accuracy of the GNN is 67.94%, while that of the kNN is 60.50%.
+Note that the training and validation accuracies reported in Table I correspond to the full set of rooms, irrespective of the number of border nodes.
 VI-C Path-Planning on uHumans2 Office Scene
-Figure 4: (a) Places subgraph of the uHumans2 office scene (52m×45m) with the assigned semantic classes (green, blue, red).
-The starting node is denoted by S, and the goal node by G; (b-d) The path of each algorithm is shown in bold black.
-Expanded nodes are depicted in regular black, while unexpanded nodes in gray: (b) Path of COA* on the places subgraph; (c) Path of HCOA* on the rooms subgraph; (d) Path of HCOA* on the places subgraph.
-[…]
-The computed path of COA* in the places layer is shown in Figure 4(4(b)).
-Figures 4(4(c))-(4(d)) illustrate the results of HCOA* in the rooms and places layers, respectively.
+Figure 5: (a) Place subgraph of the uHumans2 office (52m×45m) with its semantic classes (green, blue, red); (b-d) The starting node is denoted by S, and the goal node by G.
+The path of each algorithm is shown in black.
+Expanded nodes are depicted in dark gray, while unexpanded nodes in light gray: (b) Path of COA* on the place subgraph; (c) Path of HCOA* on the room subgraph; (d) Path of HCOA* on the place subgraph.
+In this section, we perform multiple path-planning scenarios in the office 3DSG shown in Figure 1.
+The robot must navigate from a start location to a goal while satisfying the following safety constraints, prioritized in descending order: (i) avoid computers, and (ii) avoid entering room R(0), a typically crowded area.
+To enforce these constraints, we assign semantic labels to place nodes.
+Specifically, we set $\phi_{V}^{0}(v)=3$ for all $v\in C$, where $C$ denotes the set of place nodes within a radius r=3r=3m of any computer in the scene.
+Formally, $C=\{v\in G^{0}:\|x(v)-x(o)\|_{2}\leq r,\hskip 5.0pt\forall o\in O\}$ where $x(\cdot)$ denotes the location of a place node $v$ or object $o$, and $O$ is the set of all computers in the 3DSG.
+We also assign $\phi_{V}^{0}(v)=2$ for all $v\in\text{R(0)}$, thereby penalizing passing through this room.
+The rest of the nodes have $\phi_{V}^{0}(v)=1$.
+First, we evaluate the computational efficiency of the algorithm by performing 1,000 runs with fixed start and goal.
+The starting node $v_{s}$ is located in room R(2), and the goal $v_{g}$ is in room R(1).
+Figure 5(5(a)) shows the place layer, where nodes are color-coded according to their semantic class.
+The computed path of COA* in the place layer is shown in Figure 5(5(b)).
+Figures 5(5(c))-(5(d)) illustrate the results of HCOA* in the room and place layers.
 The figures also show the expanded nodes for each algorithm.
 In particular, COA* expanded nodes in two additional rooms compared to HCOA*.
-The exact number of expanded nodes is provided in Table II, where HCOA* demonstrates a 25% reduction in expanded nodes compared to COA*.
-Table II also presents the computational time of the algorithms over 1,000 runs.
-We observe that HCOA*-MC achieves the best performance in terms of computational efficiency, reducing the execution time by 16% compared to COA*.
-However, HCOA*-GNN has the highest computational time, as GNN inference can be more time-consuming than graph search in small graphs (the entire places subgraph contains only 1,314 nodes).
-It should be noted that the first run of HCOA*-GNN takes longer, likely due to compilation and GPU/CPU warm-up overhead.
-Additionally, both GNN and MC classify R(0) as class 2 and R(2) as class 3.
-However, while the GNN assigns class 3 to R(3), MC classifies it as class 1.
-The remaining rooms are not classified, as they are not adjacent to any expanded nodes.
-Finally, all three approaches successfully compute the optimal path in the places layer, as shown in Figure 4.
+TABLE II: Path-Planning on uHumans2 Office Scene.
+Algorithm Exp. Nodes Time ($10^{-3}$s) Opt. Path
+[…]
+HCOA*-kNN 412 5.6 $\pm$ 0.2
+HCOA*-GNN 412 11.5 $\pm$ 2.5
+COA* 549 4.9 $\pm$ 0.4
+The exact number of expanded nodes is provided in Table II, where HCOA* demonstrates a 25% reduction compared to COA*.
+Table II also presents the computational time of the algorithms.
+We observe that HCOA*-MC achieves the best performance in terms of computational efficiency, reducing the execution time by 14% compared to COA*.
+However, HCOA*-GNN has the highest computational time, as GNN inference can be more time-consuming than graph search in small graphs (the entire place subgraph contains only 1,314 nodes).
+Additionally, all classification methods classify R(0) as class 2 and R(2) as class 3.
+However, while GNN and kNN assign class 3 to R(3), MC classifies it as class 1.
+Although this misclassification does not impact the final path in the current scenario, it could lead to suboptimal solutions in more complex room graphs, as the robustness of the algorithm is highly dependent on the structure of the room graph.
+Finally, all four approaches successfully compute the optimal path in the place layer, as shown in Figure 5.
 It is important to note that while the first assumption of Proposition 2 holds, the second assumption does not.
-However, this does not affect the results in this case, as the optimal path in the places layer does not require deviations into rooms outside the unique path in the rooms layer to achieve a lower-cost trajectory.
-TABLE II: Path-Planning on uHumans2 Office Scene.
-Metrics HCOA*-GNN HCOA*-MC COA*
-Expanded Nodes 412 412 549
-Time ($10^{-3}$s) 38.8 $\pm$ 47.1 4.3 $\boldsymbol{\pm}$ 2.6 5.1 $\pm$ 2.5
-Optimal Path
-VII Conclusions
-In this paper, we addressed the problem of robot navigation in 3D geometric and semantic environments by leveraging the hierarchical structure of the environment for efficient path-planning.
-We introduced Hierarchical Class-ordered A* (HCOA*), an algorithm that exploits a total order over semantic classes to guide the search process while significantly reducing computational effort.
-To classify higher-layer nodes, we proposed two methods: a Graph Neural Network-based method and a Majority-Class method.
-Through simulations on a 3D Scene Graph, we showed that HCOA* effectively reduces computational cost compared to a state-of-the-art method.
-Specifically, our results show that HCOA* finds the optimal path while reducing the number of expanded nodes by 25% and achieving a 16% reduction in computational time for a typical 3DSG with 1,314 nodes.
-Moving forward, future work will focus on applying HCOA* to larger, real-world datasets to demonstrate the full potential of the algorithm.
-Additionally, we aim to develop an online path-planning framework based on HCOA*, enabling adaptive decision-making in real-time, dynamic scenarios.
+However, this does not affect the results in this case, as the optimal path in the place layer does not require deviations into rooms outside the unique path in the room layer to achieve a lower-cost trajectory.
+To further evaluate the suboptimality of HCOA*, we conducted 500 scenarios with varied start and goal locations.
+HCOA* successfully found the optimal path in 96.95% of the cases.
+This outcome is highly dependent on the representation of the room layer; modeling a room solely by its centroid can lead to suboptimal paths, particularly in elongated or non-orthogonal rooms (e.g., R(0)) where the centroid may not accurately reflect spatial connectivity (see Figure 3).
+We also report results from the MA* implementation.
+When $\alpha=2$, performance drops to 80.38% (comparing to COA*), while $\alpha=10$ yields 99.84%.
+For larger values, accuracy approaches 100%, but for $\alpha>10^{5}$, it degrades (e.g., 95.25%), indicating numerical instability.
+These findings show that MA* requires careful tuning of $\alpha$ based on the graph size, Euclidean distances, and the number of semantic classes, highlighting the benefit of using a total order as in COA* and HCOA*.
+VI-D Path-Planning on uHumans2 Subway Scene
+Figure 6: (a) uHumans2 subway scene; (b) Place subgraph of the uHumans2 subway (88m×60m) with its semantic classes (green, blue, red); (b-d) The starting node is denoted by S, and the goal node by G.
+The path of each algorithm is shown in black.
+Expanded nodes are depicted in dark gray, while unexpanded nodes in light gray: (b) Path of COA* on the place subgraph; (c) Path of HCOA* on the room subgraph; (d) Path of HCOA* on the place subgraph.
+We performed similar simulations to those in Section VI-C on the two-floor subway 3DSG, which consists of 2,732 place nodes, shown in Figure 6(6(a)).
+The safety constraints are prioritized in descending order: (i) avoid room R(5), where stair repairs are in progress, and (ii) avoid room R(0), which is crowded.
+Accordingly, we assign semantic classes as follows: $\phi_{V}^{0}(v)=3$ for all $v\in\text{R(5)}$, $\phi_{V}^{0}(v)=2$ for all $v\in\text{R(0)}$, and $\phi_{V}^{0}(v)=1$ for all other nodes.
+To evaluate the computational efficiency of the algorithms, we performed 1,000 runs with fixed start and goal locations.
+The starting node $v_{s}$ is located in room R(10), and the goal node $v_{g}$ is in room R(4).
+Figure 6(6(b)) shows the place layer, while Figures 5(5(b))-(5(d)) depict the computed paths of COA* and HCOA*, along with the corresponding expanded nodes.
+Table III summarizes the results.
+HCOA* achieves a 59% reduction in the number of expanded nodes compared to COA*.
+In terms of computational time, HCOA*-MC and HCOA*-kNN yield reductions of 49% and 15%, respectively.
+We also assessed the suboptimality of the proposed algorithm by running 500 scenarios with varied start and goal locations.
+HCOA* successfully finds the optimal path in 70.56% of the cases.
+This decrease compared to the office environment is attributed to the increased complexity introduced by the two-floor layout and the irregular geometry of the rooms.
+For comparison, the results of MA* show that with $\alpha=2$, the performance is 72.42%, while $\alpha=10$ yields 99.16%.
+For larger values, accuracy approaches 100%, but degrades for $\alpha>10^{5}$ again indicating numerical instability.
+TABLE III: Path-Planning on uHumans2 Subway Scene.
+Algorithm Exp. Nodes Time ($10^{-3}$s) Opt. Path
+[…]
+HCOA*-kNN 1029 15.6 $\pm$ 5.1
+HCOA*-GNN 1029 36.6 $\pm$ 4.4
+COA* 2480 18.3 $\pm$ 4.5
+VII Conclusion
+In this paper, we have addressed the problem of robot navigation in 3D geometric/semantic environments by leveraging the environment’s hierarchy for efficient path-planning.
+We have introduced Hierarchical Class-ordered A* (HCOA*), an algorithm that exploits a total order over semantic classes to guide the search process while significantly reducing computational effort.
+To classify higher-layer nodes, we proposed three methods: a Majority-Class method, a k-Nearest Neighbors method, and a Graph Neural Network method.
+Through simulations on two 3D Scene Graphs, we showed that HCOA* effectively reduces computational cost compared to other state-of-the-art methods.
+Specifically, our results show that HCOA* reduces the computational time of navigation by up to 50%, while maintaining near-optimal performance.
+Future work will address the limitations of using a total order over semantic classes for navigation.
+While total ordering offers a simple and interpretable mechanism for prioritizing semantic categories, it is insufficient for capturing more complex or temporally dependent constraints (e.g., avoid kitchen after visiting bathroom).
+To handle such cases, we plan to incorporate temporal logic [22] and explore partial ordering schemes.
+These enhancements will also necessitate extending our current path-planning framework to support online replanning capabilities.
+To this end, we aim to develop an online path-planning framework based on HCOA*, enabling adaptive decision-making in real-time, dynamic scenarios.